# Bayesian Experimental Design via Contrastive Diffusions

### v1

@@ Appendix E Sequential Monte Carlo (SMC)-style resampling
In static density-based BOED, the prior $p({\boldsymbol{\theta}})$ and the likelihood $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ are available in closed-form. In the sequential experiment context, we want to generate $N$ samples ${\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{N}$ from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $M$ samples ${\boldsymbol{\theta}}^{\prime}_{1},\ldots,{\boldsymbol{\theta}}^{\prime}_{M}$ from the expected posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$. As both $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})\propto\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$ can be evaluated up to a normalizing constant, it is straightforward to extend the sampling operators of Section 4 and add a resampling step to the samples ${\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{N}$ and ${\boldsymbol{\theta}}^{\prime}_{1},\ldots,{\boldsymbol{\theta}}^{\prime}_{M}$ with weights $w_{i}$ and $w^{\prime}_{j}$:
@@
In the setting of data-based BOED, we assume access to a conditional diffusion model that allows to generate samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$. The resampling scheme proposed in (Dou and Song,, 2024) can be used as is, to improve the quality of the samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ as this is an usual conditional distribution. The resampling scheme is based on the FPS update:
$p({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(t-1)},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})\propto p_{t}({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;p({\mathb{y}}^{(t-1)}|{\boldsymbol{\theta}}^{(t-1)},{\boldsymbol{\xi}})$ (51)
where $p_{t}({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})$ is given in closed form by the unconditional diffusion model and $p({\mathb{y}}^{(t-1)}|{\boldsymbol{\theta}}^{(t-1)},{\boldsymbol{\xi}})$ is given by (46). As both these distributions are Gaussian, $p({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(t-1)},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})$ can be written in closed form and resampling weights can be written as:
$w_{i}=\frac{\tilde{w_{i}}}{\sum_{i=1}^{N}\tilde{w_{i}}}\quad\text{with}\quad\tilde{w_{i}}=p({\mathb{y}}^{(t-1)}|{\boldsymbol{\theta}}_{i}^{(t)},{\boldsymbol{\xi}})$ (52)
where $p({\mathb{y}}^{(t-1)}|{\boldsymbol{\theta}}_{i}^{(t)},{\boldsymbol{\xi}})$ is tractable (see Dou and Song, (2024) for more details).
For the expected posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})\propto\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$, update (51) takes the form:
$\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(t-1)}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$ $\propto p_{t}({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;\prod_{i=1}^{N}p({\mathb{y}}^{(t-1)}_{i}|{\boldsymbol{\theta}}^{(t-1)},{\boldsymbol{\xi}})^{\nu_{i}}$
$\propto\prod_{i=1}^{N}\left(p({\boldsymbol{\theta}}^{(t-1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;p({\mathb{y}}^{(t-1)}_{i}|{\boldsymbol{\theta}}^{(t-1)},{\boldsymbol{\xi}})\right)^{\nu_{i}}$ (53)
@@
$w^{\prime}_{j}=\frac{\tilde{w^{\prime}_{j}}}{\sum_{j=1}^{M}\tilde{w^{\prime}_{j}}}\quad\text{with}\quad\tilde{w^{\prime}_{j}}=\prod_{i=1}^{N}p({\mathb{y}}^{(t-1)}_{i}|{\boldsymbol{\theta}}^{\prime(t)}_{j},{\boldsymbol{\xi}})^{\nu_{i}}\;.$
@@ F.2.1 Source example
For CoDiff, the $\nu_{i}$’s in the expected posterior distribution were set to $\nu_{i}=\frac{1}{N}$. The current prior and posterior distributions at experimental step $k$ were initialized using respectively the prior and posterior samples at step $k-1$. Design optimization was performed using the Adam optimizer with an exponential learning rate decay schedule with initial learning rate $10^{-2}$ and decay rate $0.98$. The Langevin step-size in the DiGS method Chen et al., (2024) was set to $10^{-2}$. The joint optimization-sampling loop was run for 5000 steps.
Figure 5: Several source localisation examples. Prior (left) and expected posterior (right) samples at experiment $k$. Final ${\boldsymbol{\xi}}_{k}^{*}$ value (orange cross) at the end of the optimization sequence ${\boldsymbol{\xi}}_{0},\cdot,{\boldsymbol{\xi}}_{T}$ (blue crosses). This optimization "contrasts" the two distributions by making the expected posterior "as different as possible" from the prior.
For this example, the additional Figure 5 shows samples from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ at some intermediate step $k$, which gradually concentrates around the true sources. Samples from the expected posterior distribution are shown in comparison to illustrate its contrastive nature.
@@ F.2.2 MNIST example
For the numerical example of Section 6.3, we used the MNIST dataset (LeCun et al.,, 1998), the time varying SDE (34) with a noise schedule $(b_{max}-b_{min})/(T-t_{0})+(b_{min}-b_{max})t_{0}/(T-t_{0})$ (with $b_{max}=5$, $b_{min}=0.2,t_{0}=0,T=2$). The training of the usual score matching was done for 3000 epochs with a batch size of 256 and using Adam optimizer Kingma and Ba, (2015). We used gradient clipping and the training was done on a single A100 GPU.
@@
Update equations for the sampling operators were derived from SDE (19) for the contrastive samples of the expected posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$ and (44) for samples from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$, where ${\mathb{D}}_{k-1}$ can be added in the conditioning part without difficulty. Those updates are equivalent to (53) and (51) respectively. The resampling weights were computed as in Section E.
@@
Figures 6 and 7 show additional image reconstruction processes. The digit to be recovered is shown in the first column. The successively selected masks are shown (red line squares) in the second column with the resulting gradually discovered part of the image. The reconstruction per se can be estimated from the posterior samples shown in the last 16 columns. At each experiment, the upper sub-row shows the 16 most-likely reconstructed images, while the lower sub-row shows the 16 less-probable ones. As the number of experiments increases the posterior samples gradually concentrate on the right digit.
@@
Figure 8 then shows that design optimization is effective by showing better outcomes when masks locations are optimized (second column) than when masks are selected at random centers (third column). The highest posterior weight samples in the last 14 columns also clearly show more resemblance with the true digit in the optimized case.

### v2

@@ Appendix E Sequential Monte Carlo (SMC)-style resampling
In static density-based BOED, the prior $p({\boldsymbol{\theta}})$ and the likelihood $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ are available in closed-form. In the sequential experiment context, we want to generate $N$ samples ${\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{N}$ from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $M$ samples ${\boldsymbol{\theta}}^{\prime}_{1},\ldots,{\boldsymbol{\theta}}^{\prime}_{M}$ from the pooled posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$. As both $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})\propto\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$ can be evaluated up to a normalizing constant, it is straightforward to extend the sampling operators of Section 4 and add a resampling step to the samples ${\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{N}$ and ${\boldsymbol{\theta}}^{\prime}_{1},\ldots,{\boldsymbol{\theta}}^{\prime}_{M}$ with weights $w_{i}$ and $w^{\prime}_{j}$:
@@
In the setting of data-based BOED, we assume access to a conditional diffusion model that allows to generate samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$. The resampling scheme proposed in (Dou and Song,, 2024) can be used as is, to improve the quality of the samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ as this is a usual conditional distribution. The resampling scheme is based on the FPS update: ${\boldsymbol{\theta}}^{(t)}_{j}$ is first moved using the backward SDE into ${\boldsymbol{\theta}}^{(t+1)}_{j}$ according to $p({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(T-t-1)},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})$, which satisfies
$p({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(T-t-1)},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})\propto p_{t}({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;p({\mathb{y}}^{(T-t-1)}|{\boldsymbol{\theta}}^{(t+1)},{\boldsymbol{\xi}})$ (53)
where $p_{t}({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})$ is given in closed form by the unconditional diffusion model and $p({\mathb{y}}^{(T-t-1)}|{\boldsymbol{\theta}}^{(t+1)},{\boldsymbol{\xi}})$ is given by (47). As both these distributions are Gaussian, $p({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(T-t-1)},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})$ can be written in closed form and resampling weights can be written as:
$w_{i}=\frac{\tilde{w_{i}}}{\sum_{i=1}^{N}\tilde{w_{i}}}\quad\text{with}\quad\tilde{w_{i}}=p({\mathb{y}}^{(T-t-1)}|{\boldsymbol{\theta}}_{i}^{(t)},{\boldsymbol{\xi}})$ (54)
where $p({\mathb{y}}^{(T-t-1)}|{\boldsymbol{\theta}}_{i}^{(t)},{\boldsymbol{\xi}})$ is tractable (see Dou and Song, (2024) for more details).
For the pooled posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})\propto\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$, update (53) takes the form:
$\prod_{i=1}^{N}p({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{y}}^{(T-t-1)}_{i},{\boldsymbol{\xi}},{\mathb{D}}_{k-1})^{\nu_{i}}$ $\propto p_{t}({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;\prod_{i=1}^{N}p({\mathb{y}}^{(T-t-1)}_{i}|{\boldsymbol{\theta}}^{(t+1)},{\boldsymbol{\xi}})^{\nu_{i}}$
$\propto\prod_{i=1}^{N}\left(p({\boldsymbol{\theta}}^{(t+1)}|{\boldsymbol{\theta}}^{(t)},{\mathb{D}}_{k-1})\;p({\mathb{y}}^{(T-t-1)}_{i}|{\boldsymbol{\theta}}^{(t+1)},{\boldsymbol{\xi}})\right)^{\nu_{i}}$ (55)
@@
$w^{\prime}_{j}=\frac{\tilde{w^{\prime}_{j}}}{\sum_{j=1}^{M}\tilde{w^{\prime}_{j}}}\quad\text{with}\quad\tilde{w^{\prime}_{j}}=\prod_{i=1}^{N}p({\mathb{y}}^{(T-t-1)}_{i}|{\boldsymbol{\theta}}^{\prime(t)}_{j},{\boldsymbol{\xi}})^{\nu_{i}}\;.$
@@ F.2.1 Source example
For CoDiff, the $\nu_{i}$’s in the pooled posterior distribution were set to $\nu_{i}=\frac{1}{N}$. The current prior and posterior distributions at experimental step $k$ were initialized using respectively the prior and posterior samples at step $k-1$. Design optimization was performed using the Adam optimizer with an exponential learning rate decay schedule with initial learning rate $10^{-2}$ and decay rate $0.98$. The Langevin step-size in the DiGS method Chen et al., (2024) was set to $10^{-2}$. The joint optimization-sampling loop was run for 5000 steps. Figure 5 shows samples from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$, which gradually concentrate around the true sources as $k$ increases. The additional Figure 6 shows, at some intermediate step $k$, samples from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and from the pooled posterior distribution in comparison, to illustrate its contrastive nature.
Figure 5: Source localization example. Experiments 0 (prior samples), 4, 8 and 12. As new design locations are selected (orange crosses), samples concentrate to the true sources (red crosses). Samples with lower weights in blue, higher weights in yellow.
Figure 6: Several source localisation examples. Prior (left) and pooled posterior (right) samples at experiment $k$. Final ${\boldsymbol{\xi}}_{k}^{*}$ value (orange cross) at the end of the optimization sequence ${\boldsymbol{\xi}}_{0},\cdot,{\boldsymbol{\xi}}_{T}$ (blue crosses). This optimization "contrasts" the two distributions by making the pooled posterior "as different as possible" from the prior.
@@ F.2.2 MNIST example
In this example, the likelihood easily derives from ${\mathb{Y}}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\boldsymbol{\eta}}$, where ${\mathb{Y}},{\boldsymbol{\theta}}$ and ${\boldsymbol{\eta}}$ are familiarly seen as arrays of pixels. The transformation $T_{{\boldsymbol{\xi}},{\boldsymbol{\theta}}}$ is simply $T_{{\boldsymbol{\xi}},{\boldsymbol{\theta}}}({\mathb{U}})={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\mathb{U}}$ with ${\mathb{U}}={\boldsymbol{\eta}}$ a Gaussian variable. However, to fit in our theoretical framework, images need to be treated as mappings over a continuous spatial domain, so that ${\boldsymbol{\xi}}$ can be seen as a continuous parameter and ${\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}$ be differentiable with respect to ${\boldsymbol{\xi}}$. This is common in image processing and analysis, where 2D images are often regarded as mappings over a continuous 2D domain (see e.g. Rein van den Boomgaard and Leo Dorst, 2021 Lecture Notes) with $x_{1}$ and $x_{2}$ axes. That is, for $(x_{1},x_{2})\in\mathbb{R}^{2}$, ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$, we consider that an array of pixels ${\boldsymbol{\theta}}$ is a discrete sampled representation of a 2D function ${\boldsymbol{\theta}}(x_{1},x_{2})$ and more generally define, using abusively the same notation for continuous and sampled representations,
${\mathb{Y}}(x_{1},x_{2})={\mathb{A}}_{\boldsymbol{\xi}}({\boldsymbol{\theta}}(x_{1},x_{2}))+{\boldsymbol{\eta}}(x_{1},x_{2})$
where ${\mathb{A}}_{\boldsymbol{\xi}}(\cdot)$ is a masking operator depending on some length $h$ and defined by
${\mathb{A}}_{\boldsymbol{\xi}}({\boldsymbol{\theta}}(x_{1},x_{2}))$ $=$ ${\boldsymbol{\theta}}(x_{1},x_{2})\quad\mbox{ if }(x_{1},x_{2})\in S_{{\boldsymbol{\xi}},h}$
[…]
with $S_{{\boldsymbol{\xi}},h}=\{(x_{1},x_{2})\in\mathbb{R}^{2},\xi_{1}-h\Delta_{1}\leq x_{1}\leq\xi_{1}+h\Delta_{1},\xi_{2}-h\Delta_{2}\leq x_{2}\leq\xi_{2}+h\Delta_{2}$, denoting by $\Delta_{1}$ and $\Delta_{2}$ the sampling distances along the $x_{1}$ and $x_{2}$ axes.
To be able to derive with respect to ${\boldsymbol{\xi}}$, we then need to consider a smooth version ${\boldsymbol{\mu}}_{{\boldsymbol{\xi}},{\mathb{s}}}$ of ${\mathb{A}}_{\boldsymbol{\xi}}$. This is classically done by convoluting with a 2D Gaussian kernel $G_{\mathb{s}}$, e.g. the product of two 1D Gaussian kernels with positive scales ${\mathb{s}}=(s_{1},s_{2})$. In practice, this consists in smoothing the sharp borders of the mask, noting that ${\boldsymbol{\mu}}_{{\boldsymbol{\xi}},{\mathb{s}}}\rightarrow{\mathb{A}}_{\boldsymbol{\xi}}$ when ${\mathb{s}}\rightarrow\mathbf{0}$,
${\boldsymbol{\mu}}_{{\boldsymbol{\xi}},{\mathb{s}}}(x_{1},x_{2})$ $=$ $({\mathb{A}}_{\boldsymbol{\xi}}({\boldsymbol{\theta}})*G_{\mathb{s}})(x_{1},x_{2})$
$=$ $({\mathb{A}}_{\mathbf{0}}({\boldsymbol{\theta}})*G_{\mathb{s}})(x_{1}-\xi_{1},x_{2}-\xi_{2})$
$=$ ${\boldsymbol{\mu}}_{\mathbf{0},{\mathb{s}}}(x_{1}-\xi_{1},x_{2}-\xi_{2})$
where the second equality uses that ${\mathb{A}}_{\boldsymbol{\xi}}({\boldsymbol{\theta}}(x_{1},x_{2}))={\mathb{A}}_{\mathbf{0}}({\boldsymbol{\theta}}(x_{1}-\xi_{1},x_{2}-\xi_{2}))$. It follows that, no matter what ${\mathb{A}}_{\mathbf{0}}({\boldsymbol{\theta}})$ is, such a convolution with $G_{\mathb{s}}$ makes ${\boldsymbol{\mu}}_{\mathbf{0},{\mathb{s}}}$ into a function that is continuous and infinitely differentiable in its arguments. Then, it comes, for $i=1,2$,
$\frac{\partial{\boldsymbol{\mu}}_{{\boldsymbol{\xi}},s}}{\partial\xi_{i}}(x_{1},x_{2})$ $=$ $\frac{\partial{\boldsymbol{\mu}}_{\mathbf{0},{\mathb{s}}}}{\partial\xi_{i}}(x_{1}-\xi_{1},x_{2}-\xi_{2})$
$=$ $-\frac{\partial{\boldsymbol{\mu}}_{\mathbf{0},{\mathb{s}}}}{\partial x_{i}}(x_{1}-\xi_{1},x_{2}-\xi_{2})$
$=$ $-\left({\mathb{A}}_{\mathbf{0}}({\boldsymbol{\theta}})*\frac{\partial G_{\mathb{s}}}{\partial x_{i}}\right)(x_{1}-\xi_{1},x_{2}-\xi_{2})\;.$
These latter derivatives are all that is needed to define the gradients in our developments. If ${\mathb{A}}_{0}$ is a Heaviside step function, its convolution with a distribution leads to the cumulative density function (cdf) of this distribution. The same developments are valid by replacing the Gaussian kernel by a bivariate logistic distribution $L$ with mean $\mathbf{0}$ and scales ${\mathb{s}}=(s_{1},s_{2})$. The multivariate logistic distribution (Gumbel,, 1961; Malik and Abraham,, 1973) generalizes the univariate logistic distribution. Its pdf is $L_{2}(x_{1},x_{2};{\mathb{s}})=\frac{2!\exp(-x_{1}/s_{1}-x_{2}/s_{2})}{s_{1}s_{2}\left(1+\exp(-x_{1}/s_{1})+\exp(-x_{2}/s_{2})\right)^{3}}$. Its cdf is closed-form and is $\frac{1}{1+\exp(-x_{1}/s_{1})+\exp(-x_{2}/s_{2})}$. In practice, we simply consider a product of two independent univariate logistic distributions with mean 00 and scale $s_{i}$, $i=1,2$, $L_{1}(x_{i};s_{i})=\frac{\exp(-x_{i}/{s_{i}})}{s_{i}(1+\exp(-x_{i}/s_{i}))^{2}}$. The cdf of a 1D logistic distribution is the sigmoid function $S(x_{i};s_{i})=\frac{1}{1+\exp(-x_{i}/s_{i})}$. The convolution of a Heaviside step function with such a logistic distribution is thus a smooth sigmoid. In the MNIST example, the mask length is set to $h=7$, with sampling distances $\Delta_{1}=\Delta_{2}=1$ and we use a 2D product logistic kernel with $s_{1}=s_{2}=0.1$. Using that the 2D mask ${\mathb{A}}_{\boldsymbol{\xi}}$ can be written as the following product $\left(H(x_{1}\!-\!\xi_{1}\!+\!h)+H(\xi_{1}\!+\!h\!-\!x_{1})-1\right)\left(H(x_{2}\!-\!\xi_{2}\!+\!h)+H(\xi_{2}\!+\!h\!-\!x_{2})-1\right)$, where $H$ is the 1D Heaviside step function, it follows that the smooth ${\boldsymbol{\mu}}_{{\boldsymbol{\xi}},{\mathb{s}}}$ is
${\boldsymbol{\mu}}_{{\boldsymbol{\xi}},{\mathb{s}}}(x_{1},x_{2})\!=\!\left(S(x_{1}\!-\!\xi_{1}\!+\!h;s_{1})+S(\xi_{1}\!+\!h\!-\!x_{1};s_{1})-1\right)\left(S(x_{2}\!-\!\xi_{2}\!+\!h;s_{2})+S(\xi_{2}\!+\!h\!-\!x_{2};s_{2})-1\right).$
@@
For the numerical example of Section 6.3, we used the MNIST dataset (LeCun et al.,, 1998), the time varying SDE (34) with a noise schedule $\beta(t)=b_{min}+(b_{min}-b_{max})(t-t_{0})/(T-t_{0})$ (with $b_{max}=5$, $b_{min}=0.2,t_{0}=0,T=2$). The training of the usual score matching was done for 3000 epochs with a batch size of 256 and using Adam optimizer Kingma and Ba, (2015). We used gradient clipping and the training was done on a single A100 GPU.
@@
Update equations for the sampling operators were derived from SDE (19) for the contrastive samples of the pooled posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})$ and (C.2) for samples from the current prior $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$, where ${\mathb{D}}_{k-1}$ can be added in the conditioning part without difficulty. Those updates are equivalent to (55) and (53) respectively. The resampling weights were computed as in Section E.
@@
Figures 7 and 8 show additional image reconstruction processes. The digit to be recovered is shown in the first column. The successively selected masks are shown (red line squares) in the second column with the resulting gradually discovered part of the image. The reconstruction per se can be estimated from the posterior samples shown in the last 16 columns. At each experiment, the upper sub-row shows the 16 most-likely reconstructed images, while the lower sub-row shows the 16 less-probable ones. As the number of experiments increases the posterior samples gradually concentrate on the right digit.
@@
Figure 9 then shows that design optimization is effective by showing better outcomes when masks locations are optimized (second column) than when masks are selected at random centers (third column). The highest posterior weight samples in the last 14 columns also clearly show more resemblance with the true digit in the optimized case. The superior performance of design optimization is confirmed quantitatively in Table 1, which reports the reconstruction quality as measured by the structural similarity index measure (SSIM) (Wang et al.,, 2004), for both CoDiff and random design. 20 ground truth digit images are randomly selected and the SSIM is computed for the CoDiff and random reconstructions, after each successive experiment out of 6. Table 1 reports the median SSIM over the 20 selected digits. The SSIM is a decimal value between -1 and 1, where 1 indicates perfect similarity, 0 indicates no similarity, and -1 indicates perfect anti-correlation.